### v1

# Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning

Abstract
Reasoning is a core capability in artificial intelligence systems, for which large language models (LLMs) have recently shown remarkable progress. However, most work focuses exclusively on deductive reasoning, which is problematic since other types of reasoning are also essential in solving real-world problems, and they are less explored. This work focuses on evaluating LLMs’ inductive and abductive reasoning capabilities. We introduce a programmable and synthetic dataset, InAbHyD (pronounced in-a-bid), where each reasoning example consists of an incomplete world model and a set of observations. The task for the intelligent agent is to produce hypotheses to explain observations under the incomplete world model to solve each reasoning example. We propose a new metric to evaluate the quality of hypotheses based on Occam’s Razor. We evaluate and analyze some state-of-the-art LLMs. Our analysis shows that LLMs can perform inductive and abductive reasoning in simple scenarios, but struggle with complex world models and producing high-quality hypotheses, even with popular reasoning-enhancing techniques such as in-context learning and RLVR.  [^1] All data and generation code are freely available at https://github.com/byrantwithyou/inabhyd/.
1 Introduction
Reasoning is a hallmark of human intelligence. Peirce categorizes reasoning into deductive, inductive, and abductive reasoning  (Burch, 2024). Deductive reasoning draws a specific conclusion from a set of premises. For example, given “All humans are moral” and “Socrates is a human”, we can conclude that “Socrates is moral”. Inductive reasoning aims to identify a general principle from multiple individual cases. For example, if we see many cute cats, we may inductively reason that all cats are cute. Abductive reasoning seeks to produce hypotheses to explain observations. For example, if we hear a bird sound outside, we may abductively reason that there is a bird. The goal in both types of reasoning is to find explanations to explain observations (inductive reasoning focuses on inducing a more general rule, while abductive reasoning focuses on finding a ground fact or axiom).
The composition of inductive and abductive reasoning is ubiquitous in real-world scenarios. For example, given (i) Amy has a skin rash and leg pain and has been diagnosed with a new unknown virus. (ii) Sam has a skin rash and leg pain and has been diagnosed with the same new unknown virus. (iii) Tom also has a skin rash and leg pain. We can first induce the rule: The new unknown virus causes skin rashes and leg pain, and then we can abduce that Tom is infected with this new virus.
A fundamental difference between deductive and inductive or abductive reasoning is that deductive reasoning is deterministic. The conclusion in deductive reasoning is deterministically correct if the proof is sound and the premises are correct. However, inductive and abductive reasoning are inherently probabilistic. Technically, there may exist many conclusions, some more probable (high-quality) and some less probable (low-quality). For example, in the above bird sound case, a less probable conclusion is that someone outside is playing the bird sound on a speaker. In this regard, we view inductive and abductive reasoning as the inverse processes of deductive reasoning and study them through a unified lens. In many real-world scenarios, all three reasoning modalities are indispensable. Deductive reasoning is widely used in mathematical problem-solving, inductive reasoning is commonly used in scientific research and discovery, and abductive reasoning is often employed in medical diagnosis. Knowing and applying the principles of parsimony / Occam’s Razor is essential for LLMs to be utilized in areas such as scientific discovery, which require inductive and abductive reasoning, yet this is often overlooked. Physicists, for example, strive for the simplest explanations to account for as many observations as possible. According to our findings, LLMs struggle to generate high-quality hypotheses that match human-level intelligence. To the best of our knowledge, prior studies of inductive and abductive reasoning did not consider parsimony / Occam’s Razor in their evaluation.
The high-level research question we want to answer in this paper is: Do LLMs reason inductively and abductively? If so, how well? There is a significant body of work on evaluating the deductive reasoning capabilities of LLMs  (Saparov and He, 2023; Saparov et al., 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024), as well as techniques to improve deductive reasoning capabilities, such as in-context learning and reinforcement learning with verifiable reward, RLVR  (Lambert et al., 2024; Su et al., 2025). However, there is relatively little exploration into the inductive and abductive reasoning capabilities of large language models (LLMs). This work systematically evaluates whether and to what extent LLMs can reason abductively and inductively. We also assess whether popular reasoning-enhancing techniques, such as in-context learning  (Zhou et al., 2022; Dong et al., 2024) and RLVR  (Lambert et al., 2024; Su et al., 2025), which primarily aim to improve deductive reasoning capabilities, can also benefit inductive and abductive reasoning.
Q: All mammals are warm-blooded. Sam is a mammal. Tom is a mammal. John is a mammal. Each feline is a mammal. All canidae are mammals. All canidae are muscular. Every dog is a canidae. All wolves are canidae. All felines are not slow. Each cat is a feline. All tigers are felines. All cats are cute. Alice is a cat. All dogs are smart. Bob is a dog. All wolves are gregarious. Jessica is a wolf. All tigers are strong. Susan is a tiger. All rodents are gnawing. Each rat is a rodent. Each squirrel is a rodent. All rats are adaptable. Jack is a rat. Noah is a rat. All squirrels are bushy. Oliver is a squirrel. We observe that Sam is hairy. Alice is hairy. Bob is hairy. Fae is strong. Fae is not slow. Fae is warm-blooded. Jack is a mammal. Noah is a mammal. Oliver is a mammal.
Please produce hypotheses to explain all observations.
A: All mammals are hairy. Fae is a tiger. All rodents are mammals.
world model observations hypotheses
Figure 1: An example of a reasoning question in InAbHyD . The upper left side is the world model. Note that the dashed arrow means the subtype relationship All rats are mammals is hidden and needs to be inferred (i.e., a hypothesis to propose). Each node is a concept (e.g., cat), with its properties and members. Each edge denotes a subtype relationship (e.g., All cats are carnivores). On the upper right-hand side are observations in this reasoning example and hypotheses (ground truth answers for this reasoning example). Each hypothesis explains observations of the same color (e.g., the hypothesis All mammals are hairy explains the observations Sam is hairy; Alice is hairy; Bob is hairy). On the lower side is the natural language description of this reasoning example. Note that this example uses a real world model for illustrative purposes. In our dataset, we use fictional world models. A more detailed explanation of this example is given in Appendix A.
We create a programmable and synthetic dataset, InAbHyD (Inductive and Abductive Hypothesis Discovery), to test LLMs’ inductive and abductive reasoning capabilities. Each reasoning example has a fictional incomplete world model in the form of an ontology tree and a set of observations, represented in first-order logic. To answer each reasoning example, LLMs must produce hypotheses to explain observations under the incomplete world model. We generate synthetic examples of fictional world models to mitigate training data contamination. The ontology tree’s height characterizes the difficulty of the world model and the corresponding reasoning example. [^2] See Figure 1 for an example and Appendix A for more explanations for Figure 1 and more examples in Appendix B.
We evaluate our dataset on five widely-used LLMs: Llama3-70B  (Dubey et al., 2024) , Gemma3-27B  (Kamath et al., 2025), DeepSeek-V3  (DeepSeek-AI et al., 2024), GPT-4o  (OpenAI, 2023) , and DeepSeek-R1-Distill-Llama-70B  (DeepSeek-AI et al., 2025). For each LLM, we test whether it can produce valid and high-quality hypotheses. We also test whether popular reasoning-enhancing techniques, such as in-context learning and RLVR, which primarily were meant to improve deductive reasoning, can also benefit inductive and abductive reasoning. Our findings can be summarized as follows:
  1. 1.
    LLMs can perform inductive and abductive reasoning at high accuracy for shallow ontology trees with a single hypothesis.
  2. 2.
    The accuracy drops significantly when the ontology tree structure is complex and there are multiple hypotheses.
  3. 3.
    In some cases, LLMs may produce valid hypotheses (hypotheses which can explain all observations); but fail to produce high-quality hypotheses.
  4. 4.
    In general, we observe that reasoning-enhancing techniques for deductive reasoning only moderately benefits inductive and abductive reasoning.
2 Related Work
Evaluating deductive reasoning capabilities.
There is no consensus on whether LLMs can perform general deductive reasoning or are stochastic parrots (Bender et al., 2021). Researchers have designed different datasets and evaluation methodologies to answer this question  (Saparov et al., 2023; Saparov and He, 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024). LogiQA  (Liu et al., 2020) is a multiple-choice QA dataset to test multiple types of deductive reasoning of language models. However, it only attends to label accuracy (the correctness of the final answer) without considering the correctness of intermediate proof steps. In contrast, PrOntoQA  (Saparov and He, 2023) and PrOntoQA-OOD  (Saparov et al., 2023) evaluate both label accuracy and step-by-step proof correctness. They can synthetically generate proofs with different heights, widths, and deduction rules. Our work is similar to PrOntoQA and ProntoQA-OOD in spirit. We both use a first-order ontology tree as the world model. However, PrOntoQA and PrOntoQA-OOD do not evaluate inductive and abductive reasoning.
Evaluating non-deductive reasoning capabilities.
[…]
Dataset Controls for complexity of world models Natural language description Inductive reasoning Abductive reasoning Evaluation of hypotheses quality Multiple hypotheses
ART (Zhao et al., 2023) non-quantitative
[…]
AbductionRules (Young et al., 2022)
Wino-Why (Zhang et al., 2020)
[…]
UniADILR (Sheng et al., 2025)
Mirage (Li et al., 2025)
Case2Code (Shao et al., 2024)
InAbHyD (ours)
Table 1 summarizes the similarities and differences between our work and others on evaluating non-deductive reasoning capabilities in both the pre- and LLM eras  (Young et al., 2022; Hua et al., 2025; Zhao et al., 2023; Yang et al., 2024; Zhang et al., 2021). Our work differs from others in the following ways. First, we focus on inductive and abductive reasoning and their interactions with deductive reasoning, while other works focus only on a single reasoning modality. Second, our reasoning problem is built upon a fictional world model with natural language descriptions. The fictional world model avoids training data contamination compared with ART  (Zhao et al., 2023), where the problem is directly taken from a web corpus; the natural language descriptions of the reasoning problem render the evaluation more practical compared with symbolic reasoning, such as in InductiveBench  (Hua et al., 2025). Third, our dataset is programmable with highly-controllable difficulty parameters, allowing us to evaluate LLMs’ reasoning capabilities under various scenarios.
Techniques to enhance reasoning.
A large body of work has focused on improving the (deductive) reasoning capabilities of LLMs and related downstream tasks such as mathematics and coding  (Li, 2025; Wei et al., 2022; Yao et al., 2023; Wang et al., 2023; Su et al., 2025; Lambert et al., 2024; Zhou et al., 2022; Dong et al., 2024). The first line of work aims to elicit the emergent capabilities of LLMs by different formats of prompts, such as chain-of-thought (CoT)  (Wei et al., 2022), tree-of-thought (ToT)  (Yao et al., 2023), and self-consistency  (Wang et al., 2023). The second line of work leverages in-context learning by providing few-shot examples to LLMs. The third line of work views LLM token generation as a reinforcement learning problem and utilizes verifiable reward functions (reinforcement learning with verifiable reward, RLVRLambert et al. 2024; Su et al. 2025) to steer LLMs to improved reasoning. The fourth line of work scales LLMs’ test time to increase the accuracy of reasoning problems. It is worth noting that most techniques discussed above focus mainly on deductive reasoning. This work investigates whether the above methods benefit inductive and abductive reasoning.
3 Approach
In this section, we describe how we generate the dataset and evaluate LLMs on this dataset. At a high level, the dataset consists of synthetically-generated reasoning examples in natural language. The question description is a user prompt, and the format specification is a system prompt to LLMs. We parse LLMs’ output to evaluate their inductive and abductive reasoning capabilities on metrics such as accuracy. We acknowledge that the potential space of inductive and abductive reasoning examples is vast, so we restrict ourselves to what is expressible in first-order logic. By composing the inductive and abductive reasoning steps in Table 2, our evaluation framework can readily cover all inductive and abductive reasoning questions in first-order logic.
Each reasoning example contains a fictional world model and a set of observations. We use real person names (such as Amy) and real property names (such as strong), but we do not use real concept names (such as tiger). Instead, we use fictional concept names (such as wumpus). In doing so, LLMs will not be confused by these fictional relations, as we never generate facts that contradict those about real concepts.
LLMs must produce hypotheses to explain the observations. Our hypothesize-explain framework can cover a wide range of inductive/abductive reasoning scenarios. [^3] Traditionally, producing hypotheses to explain observations is considered abductive reasoning. However, we include inductive reasoning in the same framing. For example, given these axioms as a world model: Sally is a cat; Polly is a cat; Amy is a cat and these observations: Sally is cute; Polly is cute; Amy is cute, we can produce a hypothesis: All cats are cute to explain all observations. This is inductive reasoning. To generate a single reasoning example, we go through the following steps:
  1. 1.
    First, we construct a world model as an ontology tree in first-order logic.
  2. 2.
    Second, we traverse the ontology tree and randomly mark some axioms as hidden. Each hidden axiom is the ground truth for hypotheses.
  3. 3.
    Third, for each hidden axiom, we generate observations so that LLMs have enough information to produce ground truth hypotheses.
  4. 4.
    Fourth, we translate the incomplete world model and observations into natural language.
Subtask Reasoning type tested World model Observations Ground truth hypothesis Proof trees for observations
Infer property Inductive o(A)
o(B)
o(C)
p(A)
p(B)
p(C)
[…]
Infer
membership
relation
Abductive $\forall x(o(x)\to p_{1}(x))$
$\forall x(o(x)\to p_{2}(x))$
$\forall x(o(x)\to p_{3}(x))$
p1(A)
p2(A)
p3(A)
[…]
Infer
subtype
relation
Both o1(A)
o1(B)
o1(C)
o2(A)
o2(B)
o2(C)
[…]
Table 2: An overview of various subtasks in InAbHyD and their details in the simplest case (a height-1 ontology tree with a single hypothesis). Here o denotes an ontology concept (e.g, cat) and o(A) denotes the membership (e.g., cat(Sally) means Sally is a cat). p denotes a property (e.g, cute) and p(A) denotes A has property p (e.g., cute(Sally) means Sally is cute). We only show the proof tree for one observation for simplicity; other observations can be derived similarly. We show hypotheses green and axioms in world models purple in the proof tree.
See Table 2 for an overview of the different subtasks and their details in the simplest case (a height-1 ontology tree with a single hypothesis). We evaluate the output of LLMs (i.e., produced hypotheses) on the following metrics:
  1. 1.
    Strong accuracy: Hypotheses are strongly correct if they exactly match ground truth hypotheses.
  2. 2.
    Weak accuracy: Hypotheses are weakly correct if they explain all observations.
  3. 3.
    Hypothesis quality: A quantitative measurement of the quality of the LLMs’ produced hypotheses based on the principles of parsimony.
Next, we describe each step in more detail.
World model generation
In our approach, the world model is an ontology tree in first-order logic. Each node in the ontology tree is a concept (e.g. cat) with properties (e.g. All cats are cute, or equivalently in logical form $\forall x(\texttt{cat}(x)\rightarrow\texttt{cute}(x))$) and members (e.g. Amy is a cat, or equivalently in logical forms, cat(Amy)). Each edge in the ontology tree denotes a subtype relationship (e.g., the ragdoll node being a child of the cat node implies Each ragdoll is a cat, or equivalently in logical form, $\forall x(\texttt{ragdoll}(x)\rightarrow\texttt{cat}(x))$). We first generate the topology of the ontology tree with the height as a controllable parameter. Each non-leaf node has two or three child nodes. The height of the ontology tree is a variable that characterizes the complexity of the world model. We then make two passes of the tree in level order traversal. In the first pass, for each node, we allocate a concept name, three properties, [^4] The properties can also be negated properties, e.g., All cats are not slow. and three members. After the first pass, we have a complete world model with three types of axioms: (1) properties (e.g., All cats are cute), (2) membership relations (e.g., Amy is a cat), (3) subtype relations (e.g., Each ragdoll is a cat). In the second pass, we randomly mark some axioms as hidden. These axioms are the ground truth hypotheses. After the second pass, we have an incomplete world model.
Observation generation
Next, we generate observations so that LLMs have sufficient information to produce the ground truth hypotheses. The observations must be complex enough that the reasoning example is non-trivial, yet they should also ensure the reasoning is feasible.
  • To generate an example where the task is to infer a property (e.g., All cats are cute), we first identify the type in the ontology with this property (e.g., cat), then we sample members of this type and subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where the members have this property (e.g., Amy is cute; Sally is cute; Polly is cute).
  • To generate an example where the task is to infer a membership relation (e.g., Amy is a cat), we first identify the type in the ontology with this member (e.g., cat), then we sample properties of this type and supertypes (e.g., All cats are cute; All felines are not slow; All mammals are hairy), and add observations where this member have these properties (e.g., Amy is cute; Amy is not slow; Amy is hairy).
  • To generate an example where the task is to infer a subtype relation (e.g., All cats are felines), we first identify the subtype in the ontology (e.g., cat), then we sample members of this type and its subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where these members belong to the supertype (e.g., Amy is a feline; Polly is a feline; Sally is a feline).
Translate theories and observations into natural language
We use the same grammar as in PrOntoQA for translating logical forms into natural language. Properties are translated into the form: All cats are cute; subtype relations are randomly translated into one of three forms: Each ragdoll is a cat; Every ragdoll is a cat; All ragdolls are cats; membership relations are translated into the form Amy is a cat.
Hypothesis quality
We define a metric to evaluate the quality of hypotheses based on the principles of parsimony, or Occam’s Razor (i.e., the priciple that, other things being equal, simpler theories are better;  Baker 2022). In this regard, we encourage one hypothesis to explain as many observations as possible and penalize unnecessary hypotheses. Formally, given a world model $\mathcal{T}$, a set of observations $\mathcal{O}$, their associated proof trees $\mathcal{P}$,candidate hypotheses $\mathcal{H}$, and ground truth hypotheses $\mathcal{H}^{*}$, for $h\in\mathcal{H}$, $n(h)$ denotes the number of times $h$ appears in the proof trees of $\mathcal{O}$ (i.e., how many times $h$ is used as a premise of a proof step). We use $|\mathcal{H}|$ to denote the cardinality of $\mathcal{H}$. The quality of $\mathcal{H}$, $q(\mathcal{H})$ is defined by  [^5] We write the equation as the ratio of two averages: the numerator is the average accuracy of the predicted hypothesis (i.e., the average number of times a hypothesis appears in any valid proof), and the denominator is the average number of times a hypothesis appears in the shortest (i.e., most parsimonious) proof.:
[…]
According to the definition, the ground truth hypotheses $\mathcal{H}^{*}$ have a quality score of 1. Consider the example in Figure 1, here $\mathcal{H}^{*}$ is {Fae is a tiger, All mammals are hairy, All rodents are mammals} and $|\mathcal{H}^{*}|=3$. Each $h^{*}\in\mathcal{H}^{*}$ appears three times in the proof trees of $\mathcal{O}$, so $\sum_{h^{*}\in\mathcal{H}^{*}}n(h^{*})=9$. Consider candidate hypotheses $\mathcal{H}$ {Fae is a tiger, All mammals are hairy, All rats are mammals, All squirrels are mammals, All cats are hairy}. $|\mathcal{H}|=5$. Obviously $\mathcal{H}$ can explain $\mathcal{O}$. All rats are mammals appears twice, All squirrels are mammals appears once, and All cats are hairy appear once in the proof trees of $\mathcal{O}$, so $\sum_{h\in\mathcal{H}}n(h)=3+3+2+1+1=10$. In this case, $q(\mathcal{H})=\frac{3\times 10}{5\times 9}=0.67$.
Appendix C contains the complete algorithms for generating reasoning examples and evaluating hypotheses.
4 Results
4.1 Experimental setup
In each experiment configuration, we generate 100 independent and identically distributed reasoning examples. Each reasoning example is pure text, a user prompt with an incomplete world model and observations. We also specify the desired output format in the system prompt. To answer each reasoning example, LLMs must produce hypotheses to explain all observations. We evaluate hypotheses on three metrics: weak accuracy, strong accuracy, and quality. We compute 95% confidence intervals for both accuracy metrics using the Wilson interval (Wilson, 1927). We use the OpenAI API to evaluate GPT-4o and together.ai platform to evaluate Llama3-70B, Gemma3-27B, DeepSeek-V3, and DeepSeek-R1-Distill-Llama-70B.
In our experiments, we vary the following variables: (1) Different types of tasks (i.e., infer property, infer membership relation, or infer subtype relation), (2) Single hypothesis or multiple hypotheses, (3) The height of the ontology tree, (4) Whether the reasoning example contains in-context demonstrations and whether they are in-distribution or out-of-distribution.
4.2 Zero-shot results
To begin evaluating the inductive and abductive reasoning abilities of LLMs, we first investigate the simplest setting with zero-shot prompting.
4.2.1 Single hypothesis
We first investigate results for reasoning examples with one hypothesis for three tasks. This helps us to isolate different types of tasks and understand how LLMs perform in each case. The results are in Figure 2. In general, all models show high accuracy (both weak and strong) greater than 80% for all three tasks under a height-1 ontology tree. When the height of the ontology tree starts to increase, we also observe a general decreasing trend in all three metrics, which is not surprising. One interesting observation is that LLMs perform better in infer membership relation tasks than other tasks, especially for Gemma3-27B. For example, even under a height-4 ontology tree, Gemma3-27B can achieve near 50% weak accuracy, while the number drops to around 10% for infer subtype relation tasks. This may be because the weak correctness for infer membership relation tasks only requires searching for the concept that has some property, but do not need to understand the structure of the ontology tree.666e.g., given Amy is cute, LLMs only need to search for a concept which has the property cute to infer the membership relation of Amy. However, this is not the case for infer subtype relation and infer properties tasks. We observe a similar trend for infer subtype relation and infer properties tasks, as we can view the concept’s parent as a special property of that concept.
1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00QualityTask: infer membership relationTask: infer propertyTask: infer subtype relationGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
Figure 2: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with single hypothesis.
4.2.2 Multiple hypotheses
Next, we turn to multi-hypothesis scenarios because they are closer to real-life situations, and we are interested in understanding how LLMs perform in this setting. Each reasoning example can contain multiple hypotheses. To check the difficulty of each reasoning example, we sample the ground truth hypotheses such that the number of hypotheses grows linearly w.r.t. the height of the ontology tree. [^7] Each ground truth hypothesis has $\sim 3$ observations. So the number of observations also grows linearly w.r.t. the height of the ontology tree. Invariably, the number of axioms in the world model grows exponentially w.r.t. the height of the ontology tree. See Table 3 for statistics under multi-hypothesis scenarios and the top row of Figure 3 for results.
The height of the ontology tree # of world model axioms # of observations # of ground truth hypotheses 1 9.0 10.0 3.0 2 14.0 11.8 3.5 3 25.5 15.2 4.6 4 46.8 20.0 6.6
Table 3: Basic statistics of multi-hypothesis examples in InAbHyD . We list the average number of world model axioms, observations, and ground truth hypotheses w.r.t. the height of the ontology tree in the same set of reasoning examples we used in evaluation. Note that the number of the ground truth hypotheses and the number of observations grow linearly with the height of the ontology tree. In contrast, the number of world model axioms grows exponentially.
In general, the performance in the multi-hypothesis setting differed significantly from that in the single-hypothesis setting. For example, the accuracy of all models drops significantly when the height of the ontology tree is increased from 1 to 2, with accuracy going from above 80% to below 50%, except for GPT-4o. This is surprising because the average number of ground truth hypotheses only increases from $3.0$ to $3.5$. We attribute this to the coupling effect of multiple hypotheses to produce and the topology of the ontology tree. Reasoning examples under the ontology tree with height 3 or 4 achieve lower accuracies and quality scores for all models. We also observe a notable gap between weak accuracy and quality. For example, under the height-4 ontology tree, all models can achieve at least 20% weak accuracy, but struggle with extremely low quality and strong accuracy scores. We manually inspect some incorrect hypotheses and summarize the most common types of errors in Section 4.5.
4.3 Does in-context learning help?
[…]
Figure 3: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-short (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom). We list the absolute results of the 8-shot in-distribution and out-of-distribution in Appendix D.
LLMs have shown extraordinary generalization capabilities with few-shot examples coupled with CoT prompts (Brown et al., 2020). We are also interested in whether in-context learning can help with inductive and abductive reasoning. We add eight demonstrations to each reasoning example, with the demonstrations’ ground truth hypotheses and CoTs. The CoT consists of each observation’s step-by-step proof. We have two settings for in-context learning: (1) In-distribution demonstrations, where the demonstration has an ontology tree of the same height as the test question. (2) Out-of-distribution demonstrations, where each demonstration has a single hypothesis with a height-1 ontology tree. In both settings, we ensure demonstrations cover all types of tasks.
The in-distribution demonstrations are slightly more helpful than out-of-distribution demonstrations. This result is consistent with the hypothesis that in-context learning mimics Bayesian inference (Xie et al., 2022). We also observe that in-context learning with in-distribution demonstrations improves strong accuracy and quality in ontology trees with heights 3 and 4. We conjecture that LLMs can learn how to navigate the ontology hierarchy and produce high-quality hypotheses through in-context demonstrations. In contrast, we do not observe a significant performance improvement from out-of-distribution demonstrations.
4.4 Does RLVR help?
1234Height0.00.20.40.60.81.0Weak Accuracy1234Height0.00.20.40.60.81.0Strong Accuracy1234Height0.00.20.40.60.81.0QualityLlama3-70BDeepSeek-R1-Distill-Llama-70B
Figure 4: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree comparison results for Llama3-70B and DeepSeek-R1-Distill-Llama-70B with multiple hypotheses.
RLVR has demonstrated significant improvements in other reasoning benchmarks, and so we test whether this improvement generalizes to abductive and inductive reasoning. Specifically, we compare Llama3-70B and DeepSeek-R1-Distill-Llama-70B. In general, we observe an average performance improvement between 10% and 20%. To analyze why DeepSeek-R1-Distill-Llama-70B performs better than Llama3-70B, we also manually inspect the reasoning output of DeepSeek-R1-Distill-Llama-70B (i.e., tokens that are inside <think> and </think>). After producing hypotheses, DeepSeek-R1-Distill-Llama-70B will usually attempt to verify those hypotheses. This verification process is essentially a deductive reasoning process, and can catch some reasoning errors or fix low-quality hypotheses. A complete example is in Appendix E.
4.5 Analysis of common errors
To further understand how and why LLMs fail, we manually inspect mistaken and low-quality hypotheses. We summarize some of the most common types of errors here for further insights. We provide example responses in each category in Appendix E. (1) Wrong ontology direction.LLMs sometimes produce hypotheses in the wrong direction. For example, the ground truth hypothesis is All cats are mammals, while LLMs will produce the hypothesis All mammals are cats. (2) Ignore the ontology and produce unnecessary hypotheses.LLMs sometimes can produce unnecessary hypotheses because they ignore the ontology. For example, the ground truth hypothesis is All cats are cute, while LLMs produce two hypotheses: All cats are cute; All ragdolls are cute. (3) Fall back to trivial hypotheses.Technically, the observation is also a valid hypothesis, but it is considered low-quality because it can only explain itself. LLMs sometimes reuse observations as hypotheses. (4) Hallucinated entities.LLMs sometimes produce hypotheses with hallucinated entities (including concepts, properties, members). Usually, this happens when the entity appears in the in-context demonstrations, but LLMs use them in the test question. (5) Confuse members with concepts.We use fictional names for concepts and real-person names for members. However, LLMs sometimes do confuse these two. For example, the ground truth hypothesis is Amy is a wumpus, while LLMs produce Wumpus is Amy.
5 Conclusions and future work
In this study, we introduced a programmable and synthetic dataset, InAbHyD . Our dataset serves as an out-of-distribution evaluation of the inductive and abductive reasoning abilities of LLMs. We do not claim that success in our benchmark automatically implies success in real-world tasks, but rather, our benchmark serves as a checkpoint for developing models that are capable of robust real-world abductive and inductive reasoning.
We systematically studied LLMs’ inductive and abductive reasoning capabilities. Our results suggest that current LLMs can’t generate high-quality hypotheses that match human-level intelligence. For example, suppose we use LLMs as a scientific discovery assistant. In that case, they may generate superficial hypotheses that can explain observations but fail to uncover more general hypotheses that reveal hidden structures or connections in science. Our results can serve to better inform practitioners of LLMs of their limitations, especially in applications that rely heavily on inductive and abductive reasoning. The dataset itself opens new avenues for future work, such as training or fine-tuning models to improve their inductive and abductive reasoning capabilities.
One limitation of this work is that first-order logic (FOL) limits the expressiveness of the world model and the reasoning example. We leave it as a future work to use higher-order logic (HOL) to generate more complex world models and reasoning examples, which would require translating from HOL into natural language and vice versa. The other crucial future direction is improving the quality of LLMs’ hypotheses. For that, LLMs may benefit from RLVR  (Lambert et al., 2024) that explicitly conveys principles of parsimony in inductive and abductive reasoning with an appropriate reward. Another interesting future direction will be to determine whether large language models fine-tuned on this synthetic dataset (using supervised fine-tuning or reinforcement learning from verifiable rewards) could improve their performance on real-world tasks involving inductive and abductive reasoning. If this is the case, this synthetic dataset is more valuable because we can essentially generate an infinite number of synthetic data points, without worrying about the bottleneck of sampling data from the real world.
Appendix A Detailed explanations of Figure 1
Here we list proof trees for each observation in Figure 1. Hypotheses are in green and world model axioms are in purple.
Fae is strong
[…]
Fae is not slow
$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\neg\texttt{slow}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\neg\texttt{slow}(x))$     $\texttt{tiger}(\texttt{Fae})$                             $\neg\texttt{slow}(\texttt{Fae})$
Fae is warm-blooded
$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\texttt{mammal}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\texttt{mammal}(x))$     $\forall x(\texttt{mammal}(x)\rightarrow\texttt{warm-blooded}(x))$                              $\forall x(\texttt{tiger}(x)\to\texttt{warm-blooded}(x))$     tiger(Fae)                                                     warm-blooded(Fae)
Sam is hairy
$\forall x(\texttt{mammal}(x)\to\texttt{hairy}(x))$     mammal(Sam)           hairy(Sam)
Alice is hairy
$\forall x(\texttt{cat}(x)\to\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\to\texttt{mammal}(x))$           $\forall x(\texttt{cat}(x)\to\texttt{mammal}(x))$     cat(Alice)                           mammal(Alice)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Alice)
Bob is hairy
$\forall x(\texttt{dog}(x)\to\texttt{canidae}(x))$     $\forall x(\texttt{canidae}\to\texttt{mammal}(x))$            $\forall x(\texttt{dog}(x)\to\texttt{mammal}(x))$     dog(Bob)                             mammal(Bob)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Bob)
Jack is a mammal
rat(Jack)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Jack)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Jack)
Noah is a mammal
rat(Noah)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Noah)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Noah)
Oliver is a mammal
squirrel(Oliver)     $\forall x(\texttt{squirrel}(x)\to\texttt{rodent}(x))$             rodent(Oliver)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                                  mammal(Oliver)
Appendix B More examples
Single hypothesis; task: infer property; height-1 ontology tree.
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
Hypotheses: Dalpists are rainy.
Single hypothesis; task: infer property; height-2 ontology tree.
World Model: Rompuses are dalpists. Sarpers are dalpists. Each gergit is a dalpist. Pamela is a sarper. Jerry is a dalpist. Amy is a rompus.
Observations: Amy is rainy. Pamela is rainy. Jerry is rainy.
Hypotheses: Dalpists are rainy.
Single hypothesis; task: infer property; height-3 ontology tree.
World Model: Pamela is a porpor. Jerry is a dalpist. Twimpees are sarpers. Each gergit is a dalpist. Each storpist is a gergit. Amy is a storpist. Porpors are rompuses. Sarpers are dalpists. Rompuses are dalpists.
Observations: Jerry is not muffled. Amy is not muffled. Pamela is not muffled.
Hypotheses: Each dalpist is not muffled.
Single hypothesis; task: infer property; height-4 ontology tree.
World Model: Amy is a felper. Jerry is a dalpist. Pamela is a felper. Each rompus is a dalpist. Each porpor is a rompus. Sarpers are dalpist s. Shumples are twimpees. Gergits are dalpists. Each frompor is a porpor. Storpists are gergits. Twimpees are sarpers. Felpers are storpists.
Observations: Amy is translucent. Pamela is translucent. Jerry is translucent.
Hypotheses: Each dalpist is translucent.
Single hypothesis; task: infer membership relation; height-1 ontology tree.
World Model: Dalpists are liquid. Dalpists are rainy. Each dalpist is brown.
Observations: Jerry is brown. Jerry is rainy. Jerry is liquid.
Hypotheses: Jerry is a dalpist.
Single hypothesis; task: infer membership relation; height-2 ontology tree.
World Model: Rompuses are dalpists. Jerry is a dalpist. Each sarper is a dalpist. Each dalpist is brown. Edward is a sarper. Gergits are dalpists. Gergits are rainy. Amy is a rompus. Dalpists are liquid.
Observations: Pamela is rainy. Pamela is liquid. Pamela is brown.
Hypotheses: Pamela is a gergit.
Single hypothesis; task: infer membership relation; height-3 ontology tree.
World Model: Each sarper is a dalpist. Jerry is a dalpist. Barbara is a porpor. Each dalpist is moderate. Rompuses are dalpists. Twimpees are sarpers. Storpists are gergits. Each dalpist is large. Edward is a sarper. Pamela is a gergit. Gergits are dalpists. Debra is a storpist. Porpors are rompuses. Amy is a rompus. Twimpees are not muffled.
Observations: Nicole is moderate. Nicole is large. Nicole is not muffled.
Hypotheses: Nicole is a twimpee.
Single hypothesis; task: infer membership relation; height-4 ontology tree.
World Model: Amy is a rompus. Each porpor is a rompus. Each sarper is a dalpist. Sharon is a felper. Rompuses are dalpists. Each shumple is a twimpee. Each dalpist is fruity. Barbara is a porpor. Every frompor is a porpor. Every gergit is a dalpist. Edward is a sarper. Raymond is a frompor. Each storpist is a gergit. Jerry is a dalpist. Each dalpist is moderate. Shumples are translucent. Felpers are storpists. Nicole is a twimpee. Pamela is a gergit. Twimpees are sarpers. Debra is a storpist.
Observations: Michael is fruity. Michael is translucent. Michael is moderate.
Hypotheses: Michael is a shumple.
Single hypothesis; task: infer subtype relation; height-1 ontology tree.
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is a rompus. Jerry is a rompus. Pamela is a rompus.
Hypotheses: Dalpists are rompuses.
Single hypothesis; task: infer subtype relation; height-2 ontology tree.
World Model: Gergits are dalpists. Porpors are dalpists. Each sarper is a dalpist. Pamela is a porpor. Jerry is a dalpist. Amy is a gergit.
Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
Hypotheses: Dalpists are rompuses.
Single hypothesis; task: infer subtype relation; height-3 ontology tree.
World Model: Amy is a storpist. Jerry is a dalpist. Frompors are porpors. Each sarper is a dalpist. Each twimpee is a sarper. Pamela is a twimpee. Storpists are gergits. Porpors are dalpists. Gergits are dalpists.
Observations: Jerry is a rompus. Amy is a rompus. Pamela is a rompus.
Hypotheses: Each dalpist is a rompus.
Single hypothesis; task: infer subtype relation; height-4 ontology tree.
World Model: Each sarper is a dalpist. Jerry is a dalpist. Amy is a shumple. Pamela is a shumple. Each storpist is a gergit. Dumpuses are frompors. Gergits are dalpists. Felpers are storpists. Each twimpee is a sarper. Frompors are porpors. Shumples are twimpees. Porpors are dalpists.
Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
Hypotheses: Each dalpist is a rompus.
Multiple hypothesis; height-1 ontology tree.
World Model: Debra is a dalpist. Amy is a dalpist. Pamela is a dalpist. Edward is a dalpist. Dalpists are liquid. Each dalpist is moderate. Each dalpist is brown. Barbara is a dalpist.
Observations: Barbara is rainy. Jerry is a rompus. Jerry is liquid. Jerry is brown. Pamela is a rompus. Edward is rainy. Debra is rainy. Jerry is moderate. Amy is a rompus.
Hypotheses: Dalpists are rompuses. Dalpists are rainy. Jerry is a dalpist.
Multiple hypothesis; height-2 ontology tree.
World Model: Nicole is a gergit. Porpors are dalpists. Pamela is a sarper. Sharon is a gergit. Porpors are not cold. Dalpists are rainy. Raymond is a gergit. Each porpor is translucent. Each sarper is a dalpist. Jerry is a dalpist. Edward is a dalpist. Every gergit is a dalpist. Debra is a dalpist. Barbara is a sarper.
Observations: Barbara is liquid. Amy is translucent. Sharon is moderate. Debra is liquid. Amy is rainy. Raymond is moderate. Jerry is a rompus. Amy is not cold. Pamela is a rompus. Nicole is moderate. Edward is liquid. Amy is a rompus.
Hypotheses: Dalpists are rompuses. Dalpists are liquid. Every gergit is moderate. Amy is a porpor.
Multiple hypothesis; height-3 ontology tree.
World Model: Barbara is a sarper. Sarpers are dalpists. Debra is a shumple. Every felper is a sarper. Raymond is a porpor. Sharon is a storpist. Each storpist is a gergit. Helen is a frompor. Porpors are dalpists. Each dalpist is liquid. Amy is a sarper. Each dalpist is salty. Felpers are rainy. Every shumple is a porpor. Each twimpee is a sarper. Nicole is a gergit. Each frompor is a sarper. Jerry is a dalpist. Dalpists are not muffled. Each gergit is a dalpist. Michael is a twimpee. Sarpers are amenable.
Observations: Debra is large. Amy is a rompus. Pamela is a rompus. Edward is liquid. Pamela is rainy. Edward is not muffled. Barbara is large. Pamela is not muffled. Edward is large. Edward is salty. Jerry is a rompus. Pamela is amenable.
Hypotheses: Each dalpist is a rompus. Dalpists are large. Edward is a dalpist. Pamela is a felper.
Multiple hypothesis; height-4 ontology tree.
World Model: Every kurpor is a twimpee. Each gergit is pale. Patricia is a dumpus. Timples are frompors. Sharon is a folpee. George is an orgit. Rimpees are felpers. Stephen is a shumple. Mark is an orgit. Joseph is a porpor. Andrew is an orgit. Each tumpus is a felper. Gergits are dalpists. Felpers are not dark. Helen is a timple. Jason is a tumpus. Amy is a pergit. Rimpees are angry. Porpors are dalpists. Raymond is a twimpee. Each rimpee is moderate. Linda is a tumpus. Every shumple is a porpor. Each frompor is a sarper. Every dumpus is a storpist. Gergits are amenable. Barbara is an orgit. Rachel is a frompor. Felpers are sarpers. Jacob is a felper. Janet is a tumpus. Michael is a timple. Steven is a boompist. Edward is a dalpist. Storpists are gergits. Each boompist is a frompor. Debra is a dalpist. Nicole is a sarper. Pergits are felpers. Shirley is a timple. Jerry is a dalpist. Pame la is a kurpor. Every folpee is a frompor. Each sarper is a dalpist. Eric is a storpist. Orgits are shumples. Gergits are not o paque. Twimpees are sarpers.
Observations: Linda is brown. Helen is salty. George is translucent. Jason is brown. Raymond is large. Sharon is large. Jerry is a rompus. William is angry. Barbara is liquid. Edward is liquid. Emily is amenable. Janet is brown. Emily is not opaque. Shirley is salty. William is not dark. Pamela is a rompus. Nicole is large. William is moderate. Andrew is translucent. Michael is salty. Amy is a rompus. Debra is liquid. Emily is pale. Mark is translucent.
Hypotheses: Dalpists are rompuses. Every dalpist is liquid. Emily is a gergit. Sarpers are large. Timples are salty. William is a rimpee. Tumpuses are brown. Orgits are translucent
Appendix C Algorithms
Refer to Algorithm 1 and 2.
if task is infer property then
 select a member of the root node
 randomly select a member of the leaf nodes
 randomly select a member of nonleaf nodes
 generates the observations where selected members have the property
end if
if task is infer membership then
 select a random leaf node
 randomly select a property of the leaf node
 randomly select a property of the root node
 randomly select a property of the selected node’s non-root ancestors.
 generates the observations where this entity have all properties
end if
if task is infer subtype relation then
 select a member of the root node
 randomly select a member of the leaf nodes
 randomly select a member of nonleaf nodes
 generates the observations where selected members also belong to a supertype
end if
return all world model axioms, observations, and the ground truth hypothesis
Algorithm 1 Generate a single hypothesis
// First, build the ontology tree
// $Node()$ will initialize a node object
// Add a root node
$tree\_nodes\leftarrow[[Node()]]$
for $i$ from 1 to h1h-1 do
$current\_nodes\leftarrow[]$
for $parrent\_node\in tree\_nodes[-1]$ do
  $u\leftarrow\texttt{random}(0,1)$
  if $u>0.5$ then
   for $j$ from 1 to 2 do
    current_nodes.append(Node())current\_nodes.append(Node())
   end for
  else
   for $j$ from 1 to 3 do
    current_nodes.append(Node())current\_nodes.append(Node())
   end for
  end if
end for
tree_nodes.append(current_nodes)tree\_nodes.append(current\_nodes)
end for
for $current\_nodes\in tree\_nodes$ do
for $node\in current\_nodes$ do
  // Allocate three members and properties
  for $j$ from 1 to 3 do
   $node.members.append(RandomMember())$
   $node.properties.append(RandomProperty())$
  end for
end for
end for
if single hypothesis then
 Randomly select a task type
 generates a single hypothesis
else
 generate 1 to 3 hypothesis per tree layer
1end if
Algorithm 2 Generate a reason example
Appendix D More results
Here we list the absolute results of Figure 3 in Figure 5.
1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality8-shot, out-of-distribution demonstration8-shot, in-distribution demonstrationzero-shotGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
Figure 5: Absolute results of weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-shot (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom).
Appendix E LLM example responses
Self-verification
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
Hypotheses: Dalpists are rainy.
LLM Output: <think> … Suppose Dalpists are rainy. Since Amy and Jerry and Pamela are dalpists, they are all rainy. This hypothesis can explain all observations. </think>
Wrong ontology direction
World Model: Porpors are dalpists. Nicole is a gergit. Sarpers are dalpists. Every gergit is a dalpist. Debra is a dalpist. Each dalpist is rainy. Every porpor is not cold. Raymond is a gergit. Each porpor is translucent. Amy is a porpor. Jerry is a dalpist. Pamela is a dalpist. Sharon is a gergit. Edward is a dalpist. Michael is a sarper.
Observations: Sharon is moderate. Barbara is not cold. Debra is liquid. Barbara is rainy. Barbara is liquid. Barbara is translucent. Amy is a rompus. Raymond is moderate. Jerry is a rompus. Edward is liquid. Pamela is a rompus. Nicole is moderate.
Ground truth hypotheses: Dalpists are rompuses. Every dalpist is liquid. Each gergit is moderate. Barbara is a porpor.
LLM Output: … Rompuses are dalpists. …
Ignore the ontology and produce unnecessary hypotheses
World Model: David is a sorple. Each sorple is a dropant. Each hilper is a thorpin. Rimpees are scrompists. Brian is a sorple. Dropants are gompers. Angela is a gomper. Each thorpin is a gomper. Every scrompist is a gomper.
Observations: David is rainy. Brian is rainy. Angela is rainy.
Ground truth hypotheses: Every gomper is rainy.
LLM Output: … Every gomper is rainy. Every dropant is rainy. …
Fall back to rival hypothesee
World Model: Rorpants are dolpers. Every worple is a wumpus. Each drompee is moderate. Ashley is a rorpant. Fimples are dolpers. Ronald is a wumpus. Wumpuses are gwompants. Yompins are dolpers. Brian is a dolper. Michelle is a worple. Each dolper is a gwompant. Shirley is a rorpant. Sandra is a bempin. Linda is a fimple. Drompees are pale. Jack is a yompin. Each drompee is a bempin. Bempins are cold. Bempins are gwompants. Emma is a gwompant. Michael is a gwompant.
Observations: Shirley is a borpin. Christopher is pale. Emma is a borpin. Brian is salty. Michael is salty. Christopher is a borpin. Christopher is cold. Ashley is salty. Christopher is moderate.
Ground truth hypotheses: Each gwompant is a borpin. Each gwompant is salty. Christopher is a drompee.
LLM Output: … Shirley is a borpin. …
Hallucinated entities
World Model: Jonathan is a rifpist. Yimples are dull. Samantha is a rifpist. Karen is a gwompant. Scott is a jelgit. Olivia is a frompor. Maria is a lirpin. Twimpees are lirpins. Each rifpist is aggressive. Frompors are jelgits. Jessica is a jelgit. Amanda is a twimpee. Rifpists are jelgits. Each lirpin is a jelgit. Yimples are small. Each gwompant is a frompor. Each yimple is a rifpist.
Observations: Mary is not overcast. Scott is not overcast. Jonathan is not overcast. Samantha is a kurpor. Mary is aggressive. Jessica is a kurpor. Olivia is a kurpor. Mary is small. Mary is dull.
Ground truth hypotheses: Every jelgit is a kurpor. Jelgits are not overcast. Mary is a yimple.
LLM Output: … Mary is a dolphant. …
Confuse members with concepts
World Model: Delpees are impuses. Yumpuses are delpees. Rebecca is an impus. Kathleen is a rifpist. Each impus is rainy. Thomas is a delpee. Lompees are arpers. Each bempin is an impus. Stephen is a bempin. Every impus is shy. Laura is a yumpus. Every arper is an impus. Stephanie is an arper. Lompees are moderate. Rifpists are bempins.
Observations: Frank is moderate. Frank is rainy. Frank is shy.
Ground truth hypotheses: Frank is a lompee.
LLM Output: … Lompee is Frank. …